# Revisiting the predictability of dynamical systems: a new local data-driven approach

### v1

@@ 1 Introduction
To further show how $\alpha_{\eta}$ reveals phase-space and time features that can be used for understanding the predictability of dynamical systems, we present the evolution of the average $\alpha_{\eta}$ for seven different clusters (or regions) of the Lorenz-63 attractor. These clusters are depicted in Fig. 3, and are obtained by applying a K-means clustering method to the Lorenz-63 data used for Fig. 2, similar to the one adopted in [60]. Interestingly, the clustering identifies portions of the attractor in the phase space that seem to match some of the observations made for Fig. 2. In particular, cluster 4 represents the region where trajectories are switching lobes, clusters 1, 2, 6, and 7 represent the wings of the attractor, while clusters 3 and 5 represent intermediate regions that contain features from the five clusters just described. We note how, as expected, cluster 4 has the lowest short-term predictability among the regions considered, with a sudden drop at short forecasting horizons (yellow line). The wings of the attractor follow, with clusters 2 (dark orange) and 6 (green) being the least predictable after cluster 4, immediately followed by clusters 1 (red) and 7 (blue). Clusters 3 and 5, after a sudden predictability drop similar to the one experienced for cluster 4, maintain relatively good predictability for larger forecasting horizons compared to all other clusters. This is confirmed, if we look at the area under each curve (AUC) reported in Figure 3, that represent the average predictability of each cluster with respect to the forecasting horizons, with clusters 3 and 5 having the highest average predictability, and cluster 4 the lowest.
We also note that, despite the general trend of predictability decreasing as the prediction horizon increases, it is interesting to note that $\alpha_{\eta}$ does not always decrease monotonically. There are instances showing returns of predictability, as also reported in [61], and as reflected more clearly in Fig. 3(b); see also discussion above regarding the relationship between the definitions provided in Eqs. 8 and 9.
@@
Other than the Lorenz system, results for other idealized dynamical systems are also presented to assess the behavior of our local predictability index in notable cases. In particular, we consider the chaotic Rössler system, and the stochastic spring-slider system introduced in [16] to describe laboratory and slow earthquakes (Supplementary Information Section C). The results for all these additional systems are as expected. For the Rössler system, we obtain a similar behavior as the Lorenz-63 system. For the stochastic spring-slider system, we obtain a relatively low predictability, that is expected given the stochastic nature of the system. We note that the results included in the Supplementary Information is representative of several additional test cases that were performed as part of this study, and that can be found in the repository containing the code for the local predictability index that is provided as part of this study.
We additionally compare our results with the nonlinear local Lyapunov exponent (NLLE) [36, 37], an established index used to quantify local predictability that we find closely relevant, as reported in Supplementary Information Section D. The results for the NLLE display some notable differences with our local predictability index while also showing some similarities. These differences are arguably due to the use of different perspectives in defining predictability. Their framework stems from the Lyapunov exponent, focusing on the exponential error growth rate, while our method is rooted in the concept of recurrences in dynamical system theory. We acknowledge that both methods have their own caveats: NLLE predefines the rule of error growth, while our method neglects the information on trajectories that have left the neighborhood. To overcome this, we propose an extension of our predictability index by accounting for the information given by the distances of all the neighboring trajectories – i.e., both those that remained close to the reference state and those that departed from it, as shown in Supplementary Information Section E.
@@
Furthermore, we have compared $\alpha_{\eta}$ with the previously introduced local dynamical indices. This allows us to directly compare the relationship between predictability and local dynamical properties, rather than relying solely on prior assumptions. In Figure 4, we show scatter plots of the local predictability index versus the local dimension $d$ and the persistence $\theta^{-1}$. We note that the intuition that high dimension and low persistence means low predictability does not hold for the case of the Lorenz-63 system. Specifically, states on the two wings of the Lorenz attractor exhibit low persistence (high values of $\theta$), indicating that these states are likely to leave the neighborhood immediately [62]. However, they also demonstrate high predictability over short forecasting horizons. This is not surprising when considering the dynamics on the wings of the Lorenz attractor—although these states may leave the neighborhood quickly, their neighboring trajectories do not diverge significantly, making them predictable in short forecasting horizons.
@@
In addition to the test cases presented for idealized systems, we computed $\alpha_{\eta}$ using real-world data. In particular, focused on studying the predictability of large-scale atmospheric circulation patterns in the Europe-Atlantic sector (80°W to 50°E, 22.5°N to 70°N), using state-of-the-art ERA5 reanalysis data [63] from 1979 to 2022, with a spatial resolution of 0.25°. We selected the daily mean 500 hPa geopotential height (Z500) as the representative variable, which has been frequently used to characterize circulation dynamics in the mid-latitudes [64, 65, 66, 67, 68]. This means that we consider the Z500 map for each day as a state along the system’s trajectory in phase space.
@@
Specifically, in the Euro-Atlantic sector, the definition of weather regimes offers a direct way of classifying atmospheric states. Weather regimes are defined as recurrent and quasi-stationary large-scale circulation patterns [69, 68]. Despite being a coarse characterization of atmospheric variability, weather regimes are found to be associated with conditions that are specifically favorable for certain types of weather extremes [70], and they have exhibited varying levels of predictability in real-time forecasting [71]. In this work, we have adopted the four-regime definition widely used for scientific research [72], although various definitions for weather regimes exist [72, 73]. All extended wintertime (DJFM) Z500 patterns are classified into five different categories including NAO+, Alantic Ridge (AR), NAO-, Scandinavian Blocking (SB) and no regime, using a method slightly adapted from [74] (see Supplementary Information Section Definition of North Atlantic Wintertime Weather Regimes for details).
@@
Figure 5: Predictability analysis of North Atlantic wintertime weather regimes. (a-c) Probability distribution of $\alpha_{\eta}$ for four weather regime and no regime: NAO+ (blue dashed line), Atlantic Ridge (red dotted line), NAO- (green dot-dashed line), Scandinavian Blocking (magenta loosely dashed line), and no regime (black line), for $\eta=1-3$ days. The markers on the x-axis show the mean value of $\alpha_{\eta}$ for different categories. The statistical significance is evaluated between the distribution of each weather regime and that of no regime using the Kolmogorov-Smirnov test.
@@
Figure 5 shows the distribution of $\alpha_{\eta}$ across all five categories for $\eta=1-3$ days. The results indicate that the NAO- regime has the highest predictability, followed by NAO+, with the remaining two regimes showing no significant difference from having no regime for $\eta=1$ day. For $\eta=2$ and $3$ days, NAO+ and NAO- still maintain notably higher mean values compared to other categories, with the no regime category having the lowest values, confirming that large scale weather regimes are a primary source of predictability in the Euro-Atlantic region. These results are in excellent agreement with previous investigation dedicated to the Euro-Atlantic sector: NAO- exhibits stronger predictability than other regimes, while the blocked zonal regimes (SB and AR) are notoriously unpredictable [71, 62] [75, Section 8.2.2]). Additionally, we find further confirmation of the general fact that blockings are associated with anomalously high instability compared to zonal flow conditions [76, 77].
These findings further validate the effectiveness of our index, supported by both the underlying physical explanations and the model-based prediction results. Here we stress that we applied this index to the Euro-Atlantic region because it is an ideal test bed with well-established relevant research; however, this index can be extended to any other variables and domains.
@@
Figure 6: Comparative analysis of $\alpha_{\eta}$ with local dynamical indices for Z500 patterns in the Europe-Atlantic sector. (a-d) Scatter plots of local dimension versus local inverse persistence, where each dot represents one daily Z500 pattern. All states are colored based on $\alpha_{\eta}$ values for $\eta=1-4$ days. Note that the color scale is different in each panel in order to improve readability.
@@
Our findings demonstrate the agreement of $\alpha_{\eta}$ with known results, positioning it as a viable index for analysing the local predictability of dynamical systems. Yet, one key point remains outstanding: $\alpha_{\eta}$, as is, cannot be used for real-time applications, as it requires the future state of the system. However, since $\alpha_{\eta}$ quantifies local predictability information in the phase space, for a given unprecedented state $\boldsymbol{\zeta}^{\prime}$, we can compute the average $\alpha_{\eta}$ of its analogues (i.e., similar states to $\boldsymbol{\zeta}^{\prime}$ that appeared in the past). The number of analogues can be set once again using a quantile threshold, but this could be different from the one used to define the local properties. In practice, if we consider $N_{a}$ analogues, indicated by $\{{\bf{x}}(t_{a_{i}})\}_{i=1}^{N_{a}}$, with $t_{a_{i}}$ being the time at which the analogue number $a_{i}$ occurred, the predictability index proxy is:
@@
This can be used for real-time predictability analyses of dynamical systems. For instance, in the context of weather and climate applications, we can identify the analogues of today’s data, and compute the predictability index proxy $\hat{\alpha}_{\eta}(\boldsymbol{\zeta}^{\prime})$ using equation 10. This will give us the real-time predictability of the system as of today, thereby allowing us to adopt the index for real-time applications. An example of real-time applications is presented in Supplementary Information Section Real-time Predictability Analysis, where we show an application for the Lorenz-63 system, with the real-time predictability proxy $\hat{\alpha}_{\eta}(\boldsymbol{\zeta}^{\prime})$ successfully capturing the phase-space and time features of the predictability for the Lorenz-63 system.
@@ Conclusions
This work complements the rich body of work on the pivotal topic of predictability of dynamical systems, from the pioneering works of Poincaré and Lorenz, to modern dynamical system theory and its applications to real-life complex systems, especially in the context and weather and climate prediction. In particular, we provide a new index that can infer state-dependent predictability from high-dimensional observational datasets, overcoming the curse of dimensionality that affects more traditional approaches. While our approach has a resemblances with classical data-driven methods used for performing Lyapunov analysis on general systems [58, 78], the local predictability index is indeed novel and is able to describe flexibly key dynamical features of complex systems. We applied this index to dynamical systems of varying complexity, ranging from idealized low-dimensional systems to real-world high-dimensional atmospheric reanalysis datasets and we showcased its effectiveness. In particular, our findings confirm that when in the Euro-Atlantic sector of the mid-latitudes a blocking is present, the predictability of the atmosphere is typically anomalously low. We further showed how the new index reveals the scale-dependent nature of predictability by using different quantiles, and how it can be used in real-time problems. Our approach also addresses the ambiguity in terms of predictability that surrounds the two local dynamical indices, namely dimension $d$ and persistence $\theta^{-1}$, providing a framework that directly addresses predictability.
@@
We conclude this work with a note. The definition of $\alpha_{\eta}$, does not require $\eta>0$. Therefore, $\alpha_{\eta}$ can be computed backward in time using $\eta<0$. In this case, $\alpha_{\eta}$ characterizes the dynamics of the system as it approaches the reference state of interest, and it provides information that are entirely different from the forward-in-time analysis presented so far, using $\eta>0$. Indeed, the backward predictability analysis can be used for understanding the dynamical pathways leading to extreme events as well as other phenomena in dynamical systems [79, 80].

### v2

@@ 1 Introduction
To further show how $\alpha_{\eta}$ reveals phase-space and time features that can be used for understanding the predictability of dynamical systems, we present the evolution of the average $\alpha_{\eta}$ for seven different clusters (or regions) of the Lorenz-63 attractor. These clusters are depicted in Fig. 3, and are obtained by applying a K-means clustering method to the Lorenz-63 data used for Fig. 2, similar to the one adopted in [55]. Interestingly, the clustering identifies portions of the attractor in the phase space that seem to match some of the observations made for Fig. 2. In particular, cluster 4 represents the region where trajectories are switching lobes, clusters 1, 2, 6, and 7 represent the wings of the attractor, while clusters 3 and 5 represent intermediate regions that contain features from the five clusters just described. We note how, as expected, cluster 4 has the lowest short-term predictability among the regions considered, with a sudden drop at short forecasting horizons (yellow line). The wings of the attractor follow, with clusters 2 (dark orange) and 6 (green) being the least predictable after cluster 4, immediately followed by clusters 1 (red) and 7 (blue). Clusters 3 and 5, after a sudden predictability drop similar to the one experienced for cluster 4, maintain relatively good predictability for larger forecasting horizons compared to all other clusters. This is confirmed, if we look at the area under each curve (AUC) reported in Fig. 3, that represent the average predictability of each cluster with respect to the forecasting horizons, with clusters 3 and 5 having the highest average predictability, and cluster 4 the lowest.
We also note that, despite the general trend of predictability decreasing as the prediction horizon increases, $\alpha_{\eta}$ does not always decrease monotonically. There are instances showing returns of predictability, as also reported in [56], and as reflected more clearly in Fig. 3(b); see also discussion above regarding the relationship between the definitions provided in Eqs. 7 and 8.
@@
We additionally compare our results with the nonlinear local Lyapunov exponent (NLLE) [36, 37], an established index used to quantify local predictability that we find closely relevant, as reported in SI Appendix section 4. The results for the NLLE display some notable differences with our local predictability index while also showing some similarities (see SI Appendix Fig. S6). These differences are arguably due to the use of different perspectives in defining predictability. Their framework stems from the Lyapunov exponent, focusing on the exponential error growth rate, while our method is rooted in the concept of recurrences in dynamical system theory. We acknowledge that both methods have their own caveats: NLLE predefines the rule of error growth, while our method neglects the information on trajectories that have left the neighborhood. To overcome this, we propose an extension of our predictability index by accounting for the information given by the distance of all the neighboring trajectories – i.e., both those that remained close to the reference state and those that departed from it, as shown in SI Appendix section 5.
@@
Furthermore, we evaluated the relationship of the previously introduced local dynamical indices, $d$ and $\Theta$, to predictability, by comparing them against $\alpha_{\eta}$. In Fig. 4, we show scatter plots of the local predictability index $\alpha_{\eta}$ versus the local dimension $d$ and the inverse persistence $\theta$. We note that the intuition that high dimension and low persistence means low predictability does not hold for the case of the Lorenz-63 system. Specifically, states on the two wings of the Lorenz attractor exhibit low persistence (high values of $\theta$), indicating that these states are likely to leave the neighborhood immediately [50]. However, they also demonstrate high predictability over short forecasting horizons. This is not surprising when considering the dynamics on the wings of the Lorenz attractor—although these states may leave the neighborhood quickly, their neighboring trajectories do not diverge significantly, making them predictable in short forecasting horizons.
@@
In addition to the test case presented for idealized systems, we computed $\alpha_{\eta}$ using real-world data. In particular, we focused on studying the predictability of large-scale atmospheric circulation patterns in the Euro-Atlantic sector (see Materials and Methods for details). In this context, we consider the Z500 map for each day as a state along the system’s trajectory in the approximated phase space. We emphasize that this region serves as an ideal testbed to validate our predictability index due to its extensive body of relevant research, especially considering that no state-dependent predictability index is directly comparable for high-dimensional datasets.
@@
In the Euro-Atlantic sector, the definition of weather regimes offers a direct way of classifying atmospheric states. Weather regimes are defined as recurrent and quasi-stationary large-scale circulation patterns [57, 58]. Despite being a coarse characterization of atmospheric variability, weather regimes are found to be associated with conditions that are specifically favorable for certain types of weather extremes [59], and they have exhibited varying levels of predictability in real-time forecasting [60]. In this work, we have adopted the four-regime definition widely used for scientific research [61], although various definitions for weather regimes exist [61, 62]. Based on this definition, all extended wintertime (DJFM) Z500 daily patterns are classified into five different categories including NAO+, Alantic Ridge (AR), NAO-, Scandinavian Blocking (SB) and no regime, using a method slightly adapted from [63] (see SI Appendix section 6 for details).
@@
Figure 5: Predictability analysis of North Atlantic wintertime weather regimes. (a-c) Probability distribution of $\alpha_{\eta}$ for four weather regime and no regime: NAO+ (blue dashed line), Atlantic Ridge (red dotted line), NAO- (green dot-dashed line), Scandinavian Blocking (magenta loosely dashed line), and no regime (black line), for $\eta=1-3$ days. The markers on the x-axis show the mean value of $\alpha_{\eta}$ for different categories. The statistical significance is evaluated between the distribution of each weather regime and that of no regime using the Kolmogorov-Smirnov test. The quantile $q$ applied in this analysis is 0.99, and the Theiler window size $w$ is set to 7 days.
@@
Figure 5 shows the distribution of $\alpha_{\eta}$ across all five categories for $\eta=1-3$ days. The results indicate that the NAO- regime has the highest predictability, followed by NAO+, with the remaining two regimes showing no significant difference from having no regime for $\eta=1$ day. For $\eta=2$ and $3$ days, NAO+ and NAO- still maintain notably higher mean values compared to other categories, with the no regime category having the lowest values, confirming that large scale weather regimes are a primary source of predictability in the Euro-Atlantic region. These results are in excellent agreement with previous investigation dedicated to the Euro-Atlantic sector: NAO- exhibits stronger predictability than other regimes, while the blocked zonal regimes (SB and AR) are notoriously unpredictable [60, 50, 64]. Additionally, we find further confirmation of the general fact that blockings are associated with anomalously high instability compared to zonal flow conditions [65, 66]. These findings further validate the effectiveness of our index, supported by both the underlying physical explanations and the model-based prediction results. We note that the analysis just presented can be extended to any other variables and domains, with a key region of interest being the tropical Indo-Pacific, given the complex interaction of different modes of variability [67].
@@
Figure 6: Comparative analysis of $\alpha_{\eta}$ with local dynamical indices for Z500 patterns in the Euro-Atlantic sector. (a-d) Scatter plots of local dimension versus local inverse persistence, where each dot represents one daily Z500 pattern. All states are colored based on $\alpha_{\eta}$ values for $\eta=1-4$ days. Note that the color scale is different in each panel in order to improve readability.
@@
Other systems
In addition to the Lorenz-63 system and climate data, in Fig. 7, we demonstrate the applicability of $\alpha_{\eta}$ across a diverse range of systems. These include another canonical dynamical system (Rössler attractor), simulation data for slow earthquakes (spring-slider), a classical physics problem (double pendulum), and real-world biological data (observational ECG and EEG data). The stochastic spring-slider system was originally introduced to simulate slow earthquakes, which feature a random attractor governed by stochastic differential equations, as detailed in [16]. The final three examples use the same dataset as [68], with the double pendulum data generated through numerical simulation, while the other two are obtained from observational data.
The results for the Rössler attractor show a certain degree of similarity to those of the Lorenz-63 system: the Lorenz-63 system experiences a significant drop in predictability before lobe transitions, whereas the Rössler attractor exhibits a similar phenomenon prior to bursting events in the z-direction. We also note that the region following the bursting events exhibits markedly higher predictability, as the trajectories passing through this region will oscillate predictably in the x–y plane until re-entering the region prior to the bursting events. In the case of the spring-slider system, there is high predictability during the almost linear inter-seismic period. The time-lagged recurrence index starts decreasing when the slip rate of the slider increases. The fact that the onset of the rupture is marked by high unpredictability is in line with the idea that it is a difficult task to predict the final size of the event at the onset of the slipping phase. For the attractor of the double pendulum, we note that two endpoints of the attractor show noticeably lower predictability, whereas the transitional regions between them show relatively higher predictability, particularly for the outer parts of the attractor. In the two real-world biological data examples, we similarly observed that $\alpha_{\eta}$ reveals variability in predictability within the phase space across different forecasting horizons. These insights can improve the analysis of such systems and deepen our understanding of their underlying dynamics. The applications to these diverse systems collectively demonstrate that this novel method has broad applicability and effectiveness.
Figure 7: Summary of $\alpha_{\eta}$ applied to various systems, including another canonical dynamical system (the Rössler attractor), simulated slow earthquakes (spring-slider model), a classical physics problem (the double pendulum), and real-world biological data (observational ECG and EEG data). The forecasting horizon $\eta$ is annotated at the bottom-right corner of each attractor. All the attractors presented here are three dimensional for demonstration purposes, but their corresponding $\alpha_{\tau}$ are computed in higher-dimensional spaces, as detailed in SI Appendix section 7. The quantile $q$ applied in all analysis is 0.99, and the Theiler window size $w$ is set to 50 time steps.
@@
Our findings demonstrate the agreement of $\alpha_{\eta}$ with known results, positioning it as a viable index for analyzing the local predictability of dynamical systems. Yet, one key point remains outstanding: $\alpha_{\eta}$, as is, cannot be used for real-time applications, as it requires the future state of the system. However, since $\alpha_{\eta}$ quantifies local predictability information in the phase space, for a given unprecedented state $\boldsymbol{\zeta}^{\prime}$, we can compute the average $\alpha_{\eta}$ of its analogues (i.e., similar states to $\boldsymbol{\zeta}^{\prime}$ that appeared in the past). The number of analogues can be set once again using a quantile threshold, but this could be different from the one used to define the local properties. In practice, if we consider $N_{a}$ analogues, indicated by $\{{\bf{x}}(t_{a_{i}})\}_{i=1}^{N_{a}}$, with $t_{a_{i}}$ being the time at which the analogue number $a_{i}$ occurred, the predictability index proxy is:
@@
This can be used for real-time predictability analyses of dynamical systems. For instance, in the context of weather and climate applications, we can identify the analogues of today’s data, and compute the predictability index proxy $\hat{\alpha}_{\eta}(\boldsymbol{\zeta}^{\prime})$ using Eq. 9. This will give us the real-time predictability of the system as of today, thereby allowing us to adopt the index for real-time applications. An example of real-time applications is presented in SI Appendix section 8, where we show an application for the Lorenz-63 system, with the real-time predictability proxy $\hat{\alpha}_{\eta}(\boldsymbol{\zeta}^{\prime})$ successfully capturing the phase-space and time features of the predictability.
@@ Conclusions
This work complements the rich body of work on the pivotal topic of predictability of dynamical systems, from the pioneering works of Poincaré and Lorenz, to modern dynamical system theory and its applications to real-life complex systems, especially in the context of weather and climate prediction. In particular, we provide a time-lagged recurrence index that can infer state-dependent predictability from high-dimensional observational datasets, overcoming the curse of dimensionality that affects more traditional approaches. While our approach has a resemblances with classical data-driven methods used for performing Lyapunov analysis on general systems [53, 69], the local predictability index is indeed novel and is able to describe flexibly key dynamical features of complex systems. We applied this index to dynamical systems of varying complexity, ranging from idealized low-dimensional systems to real-world high-dimensional atmospheric reanalysis datasets and we showcased its effectiveness. In particular, our findings confirm that when in the Euro-Atlantic sector of the mid-latitudes a blocking is present, the predictability of the atmosphere is typically anomalously low. We further showed how the new index reveals the scale-dependent nature of predictability by using different quantiles, and how it can be used in real-time problems. Our approach also addresses the ambiguity in terms of predictability that surrounds the two local dynamical indices, namely dimension $d$ and persistence $\Theta$, providing a framework that directly addresses predictability.
@@
We conclude this work with a note. The definition of $\alpha_{\eta}$, does not require $\eta>0$. Therefore, $\alpha_{\eta}$ can be computed backward in time using $\eta<0$. In this case, $\alpha_{\eta}$ characterizes the dynamics of the system as it approaches the reference state of interest, and it provides information that are entirely different from the forward-in-time analysis presented so far, using $\eta>0$. Indeed, the backward predictability analysis can be used for understanding the dynamical pathways leading to extreme events as well as other phenomena in dynamical systems [70, 71]. As a methodological study, we consider such in-depth analysis to be beyond the scope of this work, but we believe it may shed new light on various scientific domains.
Materials and Methods
Local dynamical indices
In the framework proposed by Lucarini et al. [47], two dynamical indices, the local dimension $d$ and persistence $\Theta$, were introduced to characterize the local properties of dynamical systems. To compute these indices, we first consider the negative logarithmic returns $g\left({\boldsymbol{\zeta}},{\bf{x}}(t)\right)$, as defined in Eq. 1, which quantifies proximity to a given state $\boldsymbol{\zeta}$. The exceedance ${\bf{u}}({\boldsymbol{\zeta}})$, introduced in Eq. 2, is defined for the neighboring states (i.e., recurrences) of $\boldsymbol{\zeta}$ and measures how much $g\left({\boldsymbol{\zeta}},{\bf{x}}(t)\right)$ exceeds a high quantile $q$. According to extreme value theory [72, 73], assuming the independence of the exceedances, the cumulative probability distribution $F({\bf{u}},{\boldsymbol{\zeta}})$ follow the exponential member of the generalized Pareto distribution (GPD), that is:
The parameter $\sigma({\boldsymbol{\zeta}})$, which is the scale parameter of the distribution, depends on the state $\zeta$ in phase space and can be used to derive the local dimension via $d({\boldsymbol{\zeta}})=1/\sigma({\boldsymbol{\zeta}})$ [47]. In practice, the local dimension is computed by explicitly fitting an exponential distribution to the exceedances $\bf{u}({\boldsymbol{\zeta}})$, see discussion in [74] and [75]. The estimate of the attractor’s dimension derived from $d({\boldsymbol{\zeta}})$ seems to be more efficient than previous methods based on evaluating correlation integrals [76, 77]. The other local index, persistence ($\Theta$), is defined as the inverse of the extremal index $0\leq\theta\leq 1$ [78, 79], a dimensionless parameter that measures the inverse of the length of clustering of extremes. In this context, extremes are defined as the recurrences within the neighborhood around a given state in phase space. Therefore, persistence ($\Theta$) estimates the average number of consecutive recurrences, namely the time steps for the system to resemble a given state in its neighboring phase space.
This framework is the basis for the new predictability index proposed in this work, and introduced in the main text above.
500hPa Geopotential Data in the Euro-Atlantic Sector
In this study, the atmospheric circulation dynamics in the Euro-Atlantic sector (80°W to 50°E, 22.5°N to 70°N) served as a testbed to evaluate the applicability of $\alpha_{\eta}$ on a real-world high-dimensional dataset. We used the state-of-the-art ERA5 reanalysis data [80] from 1979 to 2022, with a spatial resolution of 0.25°. The daily mean 500hPa geopotential height (Z500) was selected as the representative variable, which has been frequently used to characterize atmospheric circulation dynamics in the mid-latitudes [81, 82, 83, 84, 58]. For this spatio-temporal system, we treat each daily spatial map of Z500 as an individual state, and the corresponding space can be regarded as an approximation to the phase space. This approach is widely adopted in practical applications involving real-world complex systems because accessing all variables of a system to construct its full phase space is often impossible. Instead, we can focus on a subset of available observables to approximate the system’s dynamics. In this way, the evolution of atmospheric circulation patterns can be viewed as a trajectory within this high-dimensional space, thereby enabling the application of $\alpha_{\eta}$ to assess its state-dependent predictability.
Data and Software Availability
The code for computing the local predictability index and the data used in this study are available at https://github.com/MathEXLab/TLR. The ERA5 reanalysis data used in this study are available at https://cds.climate.copernicus.eu/#!/search?text=ERA5&type=dataset. The data used for the double pendulum, ECG, and EEG are from [68].